OpenStudy (anonymous):

x^2-14x-4=0, solve by completing the square

OpenStudy (dumbcow):

here is a lesson on completing the square that might help http://www.algebra.com/my/Completing_the_Square_001.lesson?content_action=show_dev

jimthompson5910 (jim_thompson5910):

x^2-14x-4=0 x^2-14x = 4 x^2-14x + 49 = 4 + 49 ... take half of the x coefficient (-14) to get -7 and square this to get 49. Add this to both sides (x-7)^2 = 53 (x-7)^2 = 53 x - 7 = +-sqrt(53) x - 7 = sqrt(53) or x - 7 = -sqrt(53) x = 7 + sqrt(53) or x = 7 - sqrt(53) So the solutions are x = 7 + sqrt(53) or x = 7 - sqrt(53)

OpenStudy (anonymous):

x^2-14+4=0 to complete the square, we need c (ax^2+bx+c) to equal 49 since b/2 needs to equal c^2 so we will ad 53 to both sides giving us this nice equation: x^2-14x-4+53=53 x^2-14x+49=53 now we can factor easily (x-7)(x-7)=53 now put it in standard form: (x-7)^2 = 53 x-7=+or- squareroot (53) x=-squareroot(53)+7 and x=squareroot(53)+7